OpenStudy (anonymous):

how to solve arctan[tan(7pi /4)]

zepdrix (zepdrix):

\[\Large \arctan\left[\tan(7\pi/4)\right]\]So work from the inside out. Do you know what tan(7pi/4) gives us? :)

OpenStudy (ybarrap):

Recall that arctan(tan(x)) = x, because arctan is the inverse function for tan.

OpenStudy (ybarrap):

Does this make sense? What is "x" in your equation?

zepdrix (zepdrix):

@ybarrap Gotta be careful with that method since tangent is an even function! :D Luckily we were given an angle in the 4th quadrant so it will work out just fine this time.

zepdrix (zepdrix):

Just as an example: \[\Large \arctan\left[\tan(5\pi/4)\right] \ne 5\pi/4\]

zepdrix (zepdrix):

This idea applies to odd functions. Tangent is an odd function XD lol sorry typo.

OpenStudy (ybarrap):

Good point.

OpenStudy (ybarrap):

What is going on here is as follows. tan(3pi/4) is in the 2nd quadrant, so it equals \( -\sqrt{2}/2\). Since arctan(-x)=-arctan(x), we have arctan(\( -\sqrt{2}/2)=\)-arctan(\(\sqrt{2}/2)\) But arctan(\(\sqrt{2}/2)\)=\(\pi/4\), so arctan(\( -\sqrt{2}/2)=\) \(-\pi/4\), taking into account the fact that arctan(x) is odd. So the rule for arctan(x) is: arctan(tan(-x))=arctan(-tan(x))=-arctan(tan(x)), taking into account the oddness of the tan and arctan functions. For even functions, you can just do \( f^{-1}(f(x))=x\) and not worry about the sign.

OpenStudy (ybarrap):

You have \(arctan(tan(7\pi/4))\). Since \(7\pi/4\) is in the same as \(-\pi/4\) \(arctan(tan(7\pi/4))\\=arctan(tan(-\pi/4))\\=arctan(-tan(\pi/4))\\=-arctan(tan(\pi/4))\\=-\pi/4\\=7pi/4\)

OpenStudy (anonymous):

omg so much information. I am looking at 6 different charts trying to get this to make perfect sense. I understand where you get these numbers from when I look at the charts. is this mostly memorizing these charts? It is slowly making sense. Thank you for your help!! :)

OpenStudy (ybarrap):

There is logic to it all that helps to remember things. You're welcome.